OpenStudy (anonymous):

Solve the system. y = -3x + 5 and y = 2x + 5 (0, 5) (5, 0) (-5/2, 0) no solution infinitely many solutions

OpenStudy (anonymous):

(0,5)

OpenStudy (mayankdevnani):

if y = -3x + 5 and y = 2x + 5 so, \[\huge -3x+5=2x+5\] solve for x??

OpenStudy (mayankdevnani):

@charlieberzak

OpenStudy (anonymous):

yup im here

OpenStudy (mayankdevnani):

solve for x

OpenStudy (anonymous):

im thinking :P

OpenStudy (anonymous):

2x = 7x??/ something like that?

OpenStudy (mayankdevnani):

no!!

OpenStudy (anonymous):

i dont get it

OpenStudy (jhannybean):

Use the substitution method.

OpenStudy (mayankdevnani):

\[\huge -3x+5=2x+5\] \[\huge -3x-2x=5-5\] \[\huge -5x=0\] \[\huge x=0\]

OpenStudy (jhannybean):

lol.

OpenStudy (mayankdevnani):

now put x=0 in \(\large y=2x+5/-3x+5\)

OpenStudy (anonymous):

my head is starting to hurt :/ but i think im getting it

OpenStudy (jhannybean):

by the "/"he means "or"

OpenStudy (mayankdevnani):

@Jhannybean please don't give direct answers....its not our OS CoC

OpenStudy (mayankdevnani):

let @charlieberzak to solve it..only we can give hints to helper

OpenStudy (mayankdevnani):

@charlieberzak can you solve it???

OpenStudy (anonymous):

i have a similar question, i can try to solve it with ur help

OpenStudy (anonymous):

Solve the system. y = 4x + 12 and y = -2x + 6 Answer (0, 3) (-1, 8) (12, 6) no solution infinitely many solutions

OpenStudy (jhannybean):

ok, @charlie you see the y=-3x +5? and then the other equation y=2x+5? substitute the first equation for the "y"" in the second equation, can you tell me what you get?

OpenStudy (dan815):

charlie, we saw that y = those 2 equations right, y = -3x + 5 and y = 2x + 5 and -3x+5=2x+5 when we solve this we got a value for x saying at this x value these 2 equations are equal, but this doesnt exactly tell us at what y value, is it at with this x value

OpenStudy (jhannybean):

Ok.. for this one substitute the 4x+12 for the "y" in the other equation. what do you get?

OpenStudy (mayankdevnani):

@charlieberzak there is very crowd..... @Jhannybean give hint and then solve

OpenStudy (dan815):

if you think about these 2 equations y = -3x + 5 and y = 2x + 5 they are 2 equations of a line, and they will intersect at a point if they are not parallel, so we need to find which point they intersect at